# tidychangepoint: a unified framework for analyzing changepoint detection in univariate time series

### v1

@@ 3 Architecture of tidychangepoint
Every tidycpt object is created by a call to the segment() function. The method argument indicates the algorithm we want to use, and any subsequent arguments (e.g., penalty) are passed to the corresponding function (in this case, changepoint::cpt.meanvar()). Here, we pass the penalty argument to force the algorithm to use the BIC penalty function.
mlb_pelt_meanvar_bic <- mlb_hrs |>
@@ 3.2 Methods available
tidycpt objects implement methods for the generic functions as.model(), , as.segmenter(), as.ts(), changepoints(), diagnose(), fitness(), model_name(), and plot(), as well as the three generic functions from the broom package augment(), tidy(), and glance().
@@
[1] "as.ts"        "changepoints" "model_name"   "nobs"
@@
  • augment() returns a tsibble (E. Wang, Cook, and Hyndman 2020) that is grouped according to the regions defined by the changepoint set.
@@
1 changep~ 2.2.4   PELT      <list [1]> meanvar    BIC        -999. 0.018 secs
@@ 4 Segmenters
Note that while tidychangepoint uses only S3 classes, segmenters (such as those of class cpt) may belong to S4 classes. [This is one reason why as.seg_cpt(), which coverts a segmenter of any class to a mod_cpt object, is necessary. ] fitness() returns a named vector with the type and value of the penalized objective function used by the segmenting algorithm.
@@ 4.1 Segmenting algorithms provided
Currently, segment() wraps the following algorithms:
  • PELT and Binary Segmentation as implemented by changepoint (Killick and Haynes 2022)
  • Wild Binary Segmentation as implemented by wbs (Baranowski and Fryzlewicz 2019)
@@ 4.3 Spotlight: Coen’s algorithm
The goal of Coen’s algorithm (Suárez-Sierra, Coen, and Taimal 2023) is to find the optimal set of changepoints for the NHPP model described in Section 5.3. That is, given a time series $y$, find the changepoint set $\tau$ that minimizes the value of $BMDL(\tau,NHPP(y|\hat{\theta}_{\tau}))$. To find good candidates, this genetic algorithm starts with a randomly selected set of popSize changepoint sets, and then iteratively “mates” the best (according to the BMDL penalty function, see Section 6.3) pairs of these changepoint sets to recover a new “generation” of popSize changepoint sets. This process is repeated until a stopping criteria is met (most simply, maxiter times). Thus, a constant number (popSize $\times$ maxiter) possible changepoint sets are considered, and the one with the lowest BMDL score is selected.
@@
By default, the function GA::gabin_Population() selects candidate changepoints uniformly at random with probability 0.5, leading to candidate changepoint sets of size $n/2$, on average. These candidate changepoint sets are usually poor (since $n/2$ changepoints is ludicrous), and we observe much better and faster performance by seeding the first generation with smaller candidate changepoint sets. To this end, the build_gabin_population() function runs several fast algorithms (i.e., PELT, Wild Binary Segmentation, etc.) and sets the initial probability of being selected to three times the average size of the changepoint set found by these algorithms. This results in much more rapid covergence around the optimal changepoint set. Alternatively, tidychangepoint also provides the log_gabin_population() function, which sets the initial probability to $\ln{n}/n$.
@@ 5 Models
All model objects are created by calls to one of the model-fitting functions listed in Section 5.1, whose name begins with fit_. These functions all inherit from the class fun_cpt. All model objects inherit from the mod_cpt base class.
@@
 \$ data         : Time-Series [1:98] from 1 to 98: 0.002387 0.000456 0.001034 0.002604 0.003096 ...
 \$ tau          : int [1:7] 6 22 28 31 34 54 64
 \$ region_params: tibble [8 x 3] (S3: tbl_df/tbl/data.frame)
  ..\$ region           : chr [1:8] "[0,6)" "[6,22)" "[22,28)" "[28,31)" ...
  ..\$ param_mu         : num [1:8] 0.00192 -0.00104 0.00266 0.00519 0.00447 ...
  ..\$ param_sigma_hatsq: Named num [1:8] 9.99e-07 6.17e-06 4.40e-06 3.01e-06 2.58e-06 ...
  .. ..- attr(*, "names")= chr [1:8] "[0,6)" "[6,22)" "[22,28)" "[28,31)" ...
@@
 [7] "plot"      "print"     "residuals" "tidy"
@@
1 tidy~ 0.0.1   meanvar   <NULL>        7 0.00186   477. -908. -849. -877. -863.
(a) The diagnose() function applied to a model object of class mod_cpt.
(b) The diagnose() function applied to a model object of class nhpp.
Figure 4: Output from the diagnose() function.
@@ 5.1 Model-fitting functions provided
The following model-fitting functions are provided by tidychangepoint:
  • fit_meanshift_norm(): fits the meanshift model assuming a normal distribution and white noise errors with the $m+2$ parameters $\theta=\{\sigma^{2},\mu_{0},\ldots,\mu_{m}\}$.
  • fit_meanshift_lnorm(): fits the meanshift model assuming a log-normal distribution and white noise errors with the $m+2$ parameters $\theta=\{\sigma^{2},\mu_{0},\ldots,\mu_{m}\}$.
  • fit_meanshift_pois(): fits the meanshift model assuming a Poisson distribution and white noise errors with the $m+2$ parameters $\theta=\{\sigma^{2},\mu_{0},\ldots,\mu_{m}\}$.
  • fit_meanshift_norm_ar1(): fits the meanshift model assuming a normal distribution and autocorrelated AR(1) lagged errors (as decribed in Shi et al. (2022)) with the $m+3$ parameters $\theta=\{\sigma^{2},\phi,\mu_{0},\ldots,\mu_{m}\}$, where $\phi$ is the measure of autocorrelation.
  • fit_trendshift(): fits the trendshift model assuming a normal distribution and white noise errors with the $32m+$ 3 parameters $\theta=\{\sigma^{2},\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}\}$. This model is a special case of fit_lmshift().
  • fit_trendshift_ar1(): fits the trendshift model assuming a normal distribution and autocorrelated errors with the $42m+$ 4 parameters $\theta=\{\sigma^{2},\phi,\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}\}$.
  • fit_lmshift(): fits a polynomial of degree $p$ model assuming a normal distribution and white noise errors with the $p(m+1)+1$ parameters $\theta=\{\sigma^{2},\beta_{00},\ldots,\beta_{0m},\beta_{10},\ldots,\beta_{pm}\}$. This model uses the lm() function from the stats, which provides flexibility.
  • fit_lmshift_ar1(): fits a polynomial of degree $p$ model assuming a normal distribution and autocorrelated errors with the $p(m+1)+1$ parameters $\theta=\{\sigma^{2},\beta_{00},\ldots,\beta_{0m},\beta_{10},\ldots,\beta_{pm}\}$.
  • fit_meanvar(): fits the meanvar model with the $22m+$ 2 parameters $\theta=\{\sigma_{0}^{2},\ldots,\sigma_{m}^{2},\mu_{0},\ldots,\mu_{m}\}$.
  • fit_nhpp(): fits the non-homogeneous Poisson process model described in Section 5.3 with the $22m+$ 2 parameters $\theta=\{\alpha_{0},\ldots,\alpha_{m},\beta_{0},\ldots,\beta_{m}\}$
@@ 5.3 Spotlight: Non-homogenous Poisson process model
Let $z(y)\subseteq\{t\}$ be a subset of the indices of the original times series $\{y_{t}\}$, for which the observation $y_{t}$ exceeds some threshold, which by default is the empirical mean $\bar{y}$. More formally, a time index $t\in z$ if and only if $y_{t}>\bar{y}$. This new time series $z(y)$ is called the exceedances of the original time series $y$. Following Suárez-Sierra, Coen, and Taimal (2023), we model the time series $z(y)$ as a non-homogenous Poisson process (NHPP), using a Weibull distribution parameterized by shape ($\alpha$) and scale ($\beta$). In a Bayesian setting, those parameters follow a Gamma distribution with their appropriate hyperparameters. Thus, we assume that there exists a set of parameters $\theta=(\alpha_{0}\ldots\alpha_{m},\beta_{0},\ldots,\beta_{m})$ for which the exceedances $z(y)$ of the original time series $y$ are modeled accurately as a non-homogenous Poisson process.
The best-fit parameters are determined by maximizing the log posterior function described below
$\ln{\left(g(\theta_{\tau})|y\right)}\propto\ln{L_{NHPP}(y|\theta_{\tau})}+\ln{g(\theta_{\tau})}\,.$
@@ 6 Penalized objective functions
We call the function $f$ a penalized objective function because it comes in the form
$f(\tau,M(y|\theta_{\tau}))=P_{f}(\tau)-2\ln{L_{M}(y|\theta_{\tau})}\,,$
where $P_{f}$ is a function that guards against overfitting by penalizing large changepoint sets—that is, it is a monotonically increasing function of $m$. [Occasionally, we may abuse terminology by referring to $f$ as a penalty function.] The stats package (R Core Team 2024) provides generic methods for the well-known penalty functions AIC() and BIC(), which work seamlessly on tidychangepoint models because they all implement methods for logLik(). In addition, tidychangepoint provides generic functions and methods for three additional penalty functions: MBIC(), MDL(), and BMDL().
@@
6.1 Modified Bayesian Information Criterion
@@
$P_{MBIC}(\tau)=3m\ln{n}+\sum_{j=1}^{m+1}\ln{\frac{\ell_{j}}{n}}\,.$
6.2 Minimum Descriptive Length
@@
$P_{MDL}(\tau)=\frac{a(\theta_{\tau})}{2}\cdot\sum_{j=0}^{m}\log{\ell_{j}}+2\ln{m}+\sum_{j=2}^{m}\ln{\tau_{j}}+\left(2+b(\theta_{\tau})\right)\ln{n}\,.$
@@
6.3 Bayesian Minimum Descriptive Length
@@
$BMDL(\tau,NHPP(y|\hat{\theta}_{\tau})=P_{MDL}(\tau)-2\ln{L_{NHPP}(y|\hat{\theta}_{\tau})}-2\ln{g(\hat{\theta}_{\tau})}\,.$
@@ 6.4 Extending tidychangepoint to include new penalty functions
 - attr(*, "df")= num 23
 - attr(*, "nobs")= int 98
 - attr(*, "tau")= int [1:7] 6 22 28 31 34 54 64
@@ 7 Results
In this section we illustrate how tidychangepoint can be useful in determining and analyzing possible sets of changepoints on both simulated and real data.
7.1 Recap of MLB example
In Section 2 and Section 3, we computed created five different tidycpt objects, each of which contains information about a different combination of algorithm, model, and penalty function.
mods <- list(
  mlb_pelt, mlb_pelt_meanshift, mlb_pelt_meanvar_bic, mlb_ga_shi, mlb_coen
Using map() (from the purrr package (Wickham and Henry 2023)) and glance(), we can easily compare our results, although care must be taken to ensure that those comparisons are meaningful. In this case, we build a list object of the various tidycpt objects, then iterate the glance() function over that list with the map() function. Since glance() always returns a one-row tibble, we can then combine those tibbles into one using the list_rbind() function. While comparing, say, the elapsed time across these algorithms and models is fair, comparing the fitness values is more fraught. BMDL scores are not comparable to BIC or MBIC scores, for example.
mods |>
  map(glance) |>
  list_rbind()
# A tibble: 5 x 8
  <chr>    <pckg_> <chr>     <list>     <chr>      <chr>      <dbl> <drtn>
1 changep~ 2.2.4   PELT      <list [1]> meanvar    MBIC       -920.   0.077 secs
2 changep~ 2.2.4   PELT      <list [1]> meanshift~ MBIC       -852.   0.203 secs
3 changep~ 2.2.4   PELT      <list [1]> meanvar    BIC        -999.   0.018 secs
4 GA       3.2.4   Genetic   <list [1]> meanshift~ BIC        -932. 120.480 secs
5 GA       3.2.4   Genetic   <list [1]> nhpp       BMDL        153.  63.060 secs
Consider the distribution of the changepoints identified by our five algorithms.
mods |>
  map(changepoints) |>
  list_c() |>
  table()
 6  7 22 23 28 31 34 35 36 54 64
 1  1  2  1  1  1  1  1  1  1  1
We see that a changepoint at 22 or 23 occurs three times, while a changepoint at 34, 35, or 36 also occurs three times. We also have two votes for a changepoint at 6 or 7. Let us create a vector of candidate changepoints with these values, which we might consider as a sort of algorithmic consensus. These correspond to the years show below, which are in the ballpark with the changepoints suggested by Woltring, Rost, and Jubenville (2018) and B. James (2014), but far from the true changepoints of 1973 and 2022222Recall, however, that since 2022 is an endpoint, it is not allowed to be a changepoint..
cpt_algs <- c(72235)
mlb_pelt\$time_index[cpt_algs] |>
  as_year()
[1] "1931" "1946" "1959"
cpt_true <- c(49)
mlb_pelt\$time_index[cpt_true] |>
  as_year()
[1] "1973"
The PELT algorithm, using the meanvar model and the MBIC criteria, did return a subset of cpt_algs (namely, the years 1946, and 1959). But why weren’t the candidate changepoint sets cpt_algs or cpt_true identified by our algorithms? We can use tidychangepoint to dig deeper.
Our target is the value of the penalized objective function MBIC() returned by PELT after searching with the meanvar model. The fitness() function gives us this value.
fitness(mlb_pelt)
     MBIC
-920.4666
Because PELT is known to be optimal under mild conditions, all other candidate changepoints sets should return a value that is worse. [Note that we are in the unusual but valid situation in which the log likelihood is positive!] As expected, both of these MBIC values are worse (further from $\infty$- ∞).
list(cpt_algs, cpt_true) |>
  map(fit_meanvar, x = mlb_hrs) |>
  map_dbl(MBIC)
[1] -905.7639 -881.1987
It must be the case that adding the changepoint 1931 to the set that already included 1946 and 1959 resulted in a worse MBIC score. Moreover, the true changepoint set of 1973 was even worse.
Similarly, we can compare these candidate changepoint sets to the results from Coen’s algorithm, which uses the NHPP model and the BMDL penalty. However, this algorithm found no changepoints at all. Note that Coen’s algorithm is only a heuristic, and so unlike in the case of PELT, it is possible that we could manually identify a changepoint set with a lower BMDL than the one found by the algorithm. However, in this case, we don’t.
fitness(mlb_coen)
   BMDL
153.082
list(cpt_algs, cpt_true) |>
  map(fit_nhpp, x = mlb_hrs) |>
  map_dbl(BMDL)
[1] 209.5659 182.2870
7.2 Particular matter in Bogotá, Colombia
We can use the PELT algorithm as implemented in the changepoint package through the tidychangepoint interface by calling the segment() function and specifying pelt as the method argument.
bog_pelt <- segment(bogota_pm, method = "pelt")
With the default options PELT performs curiously on these data, identifying pairs of nearly consecutive changepoints at two points in time (12-14 and 1023-1025).
changepoints(bog_pelt)
[1]   12   14  393  468  821 1023 1025
However, using the meanvar model instead of the meanshift model improves the situation considerably. Note that we have set the minseglen argument to 3 to avoid unreasonably short regions.
bog_pelt_meanvar <- bogota_pm |>
  segment(method = "pelt"model_fn = fit_meanvar, minseglen = 3)
length(changepoints(bog_pelt_meanvar))
[1] 4
The diagnose() function produces the informative data graphic shown in Figure 6. The figure shows the original time series (black line) segmented by the changepoints determined by the PELT algorithm (vertical dotted lines). The horizontal red lines indicate the mean and 1.96 times the standard deviation within each region, as per the summary produced by tidy(). The lower plot shows the distribution of the residuals.
Figure 6: Diagnostic plot for the PELT algorithm applied to the Bogotá particulate matter data.
A careful look at the data suggest that the trendshift model, which allows for a linear slope within each region, might be more appropriate. Setting the method argument to segment() to ga allows us to employ a genetic algorithm for which we can specify this model via the model_fn argument. Here, we also use the MDL() penalty function and the log_gabin_population() function.
bog_ga <- bogota_pm |>
  segment(
    method = "ga"model_fn = fit_trendshift, penalty_fn = MDL,
    population = log_gabin_population(bogota_pm),
    maxiter = 500run = 100
In this case, the algorithm finds 4 changepoints.
changepoints(bog_ga, use_labels = TRUE)
[1] "2018-11-06" "2019-04-19" "2019-12-20" "2020-03-29"
Note that the MDL score reported by the genetic algorithm is better (lower) than that of the corresponding changepoint set returned by PELT. This may not be surprising, given that PELT minimized the MBIC penalty value fit to a different model, but nevertheless it suggests that genetic algorithms have something to contribute.
fitness(bog_ga)
     MDL
8803.814
bog_pelt |>
  changepoints() |>
  fit_trendshift(x = bogota_pm) |>
  MDL()
[1] 8843.263
bogota_pm |>
  fit_trendshift_ar1(tau = changepoints(bog_ga)) |>
  MDL()
[1] 8770.549
Figure 7 shows that the trendshift model fits the data very well.
Figure 7: Diagnostic plot for a genetic algorithm with a trendshift model applied to the Bogotá particulate matter data.

### v2

@@ 3 Architecture of tidychangepoint
Every tidycpt object is created by a call to the segment() function. As a verb, “segment” describes the action of splitting a time series into regions, and its usage accords with the tidyverse convention of using verbs as function names. The method argument indicates the algorithm we want to use, and any subsequent arguments (e.g., penalty) are passed to the corresponding function. Here, we pass the penalty argument to force changepoint::cpt.meanvar() to use the BIC penalty function.
mlb_pelt_meanvar_bic <- mlb_bavg |>
@@ 3.2 Methods available
tidycpt objects implement methods for the generic functions as.model(), as.segmenter(), as.ts(), changepoints(), diagnose(), fitness(), model_name(), regions, and plot(), as well as three generic functions from the broom package: augment(), tidy(), and glance().
@@
[1] "as.ts"        "changepoints" "model_name"   "nobs"         "plot"
[6] "print"
@@
  • augment() returns a tsibble (E. Wang, Cook, and Hyndman 2020) that is grouped according to the regions defined by the changepoint set. This object class is appropriate for this use case, since the grouping by changepoint region is natural in this context, and a tsibble is a grouped tibble that is aware of its indexing variable. Users can of course undo the grouping with dplyr::ungroup() or coerce to a regular tibble with tibble::as_tibble().
@@
1 changep~ 2.3     PELT      <list [1]> meanvar    BIC        -782. 0.009 secs
@@ 4 Segmenters
Note that while tidychangepoint uses only S3 classes, segmenters (such as those of class cpt) may belong to S4 classes. This is one reason why as.seg_cpt(), which coverts a segmenter of any class to a mod_cpt object, is necessary. fitness() returns a named vector with the type and value of the penalized objective function used by the segmenting algorithm.
@@ 4.1 Segmenting algorithms provided
Table LABEL:tbl-coverage shows the values passable to the method argument of the segment() function and what they do. Currently, segment() wraps the following algorithms:
  • PELT and Binary Segmentation as implemented by changepoint (Killick 2024)
  • Wild Binary Segmentation as implemented by wbs (Baranowski and Fryzlewicz 2024)
@@
Table 1: Segmenting methods provided by tidychangepoint.
Method Pkg Segmenter Class Helper Wraps
binseg changepoint cpt NA changepoint::cpt.meanvar()
coen tidychangepoint seg_basket segment_coen() NA
ga GA tidyga segment_ga() GA::ga()
ga-coen GA tidyga segment_ga_coen() segment_ga()
ga-shi GA tidyga segment_ga_shi() segment_ga()
manual tidychangepoint seg_cpt segment_manual() NA
null tidychangepoint seg_cpt segment_manual() NA
pelt changepoint cpt segment_pelt() changepoint::cpt.mean() or changepoint::cpt.meanvar()
random GA tidyga segment_ga_random() segment_ga()
segneigh changepoint cpt NA changepoint::cpt.meanvar()
single-best changepoint cpt NA changepoint::cpt.meanvar()
wbs wbs wbs NA wbs::wbs()
@@ 4.3 Spotlight: Coen’s algorithm
The goal of Coen’s algorithm (Suárez-Sierra, Coen, and Taimal 2023) is to find the optimal set of changepoints for the NHPP model described in Section 5.3. That is, given a time series $y$, find the changepoint set $\tau$ that minimizes the value of $BMDL(\tau,NHPP(y|\hat{\theta}_{\tau}))$. To find good candidates, this genetic algorithm starts with a randomly selected set of popSize changepoint sets, and then iteratively “mates” the best (according to the BMDL penalty function, see Section 6.1.4) pairs of these changepoint sets to recover a new “generation” of popSize changepoint sets. This process is repeated until a stopping criteria is met (most simply, maxiter times). Thus, a constant number (popSize $\times$ maxiter) possible changepoint sets are considered, and the one with the lowest BMDL score is selected.
@@
By default, the function GA::gabin_Population() selects candidate changepoints uniformly at random with probability 0.5, leading to candidate changepoint sets of size $n/2$, on average. These candidate changepoint sets are usually poor (since $n/2$ changepoints is ludicrous), and we observe much better and faster performance by seeding the first generation with smaller candidate changepoint sets. To this end, the build_gabin_population() function runs several fast algorithms (i.e., PELT, Wild Binary Segmentation, etc.) and sets the initial probability of being selected to three times the average size of the changepoint set found by these algorithms. This results in much more rapid covergence around the optimal changepoint set. Alternatively, tidychangepoint also provides the log_gabin_population() function, which sets the initial probability to $\ln{n}/n$ (see Section 7.2).
@@ 5 Models
All model objects are created by calls to one of the model-fitting functions listed in Section 5.1, whose name begins with fit_. These functions all inherit from the class fun_cpt and can be listed using ls_models(). All model objects inherit from the mod_cpt base class.
@@
 \$ data         : Time-Series [1:95] from 1 to 95: -0.000173 -0.001663 -0.003656 0.000285 0.010521 ...
 \$ tau          : int [1:4] 15 18 52 87
 \$ region_params: tibble [5 x 3] (S3: tbl_df/tbl/data.frame)
  ..\$ region           : chr [1:5] "[1,15)" "[15,18)" "[18,52)" "[52,87)" ...
  ..\$ param_mu         : num [1:5] -0.00173 -0.0075 0.00335 -0.00719 -0.0029
  ..\$ param_sigma_hatsq: Named num [1:5] 5.63e-05 5.07e-07 3.83e-05 9.67e-06 1.62e-06
  .. ..- attr(*, "names")= chr [1:5] "[1,15)" "[15,18)" "[18,52)" "[52,87)" ...
@@
[7] "regions"   "residuals" "tidy"
@@
1 tidy~ 1.0.0.~ meanvar   <NULL>        4 0.00507   367. -706. -671. -689. -674.
@@ 5.1 Model-fitting functions provided
Table LABEL:tbl-models lists the model-fitting functions provided by tidychangepoint, along with their parameters. Any of these functions can be passed to the model_fn argument of segment_ga() (see Section 4.1).
Table 2: Model-fitting functions provided by tidychangepoint. Unless otherwise noted, all models assume a normal distribution and white noise errors. The suffix _ar1 indicates autocorrelated AR(1) lagged errors (as decribed in Shi et al. (2022)), where $\phi$ is the measure of autocorrelation. See Section 5.3 for a description of the non-homogeneous Poisson process model used by fit_nhpp().
Function name $k$ Parameters ($\theta$) Note
fit_meanshift_norm() $m+2$ $\sigma^{2},\mu_{0},\ldots,\mu_{m}$ Normal
fit_meanshift_lnorm() $m+2$ $\sigma^{2},\mu_{0},\ldots,\mu_{m}$ log-normal
fit_meanshift_pois() $m+2$ $\sigma^{2},\mu_{0},\ldots,\mu_{m}$ Poisson
fit_meanshift_norm_ar1() $m+3$ $\sigma^{2},\phi,\mu_{0},\ldots,\mu_{m}$
fit_trendshift() $32m+$ 3 $\sigma^{2},\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}$ special case of fit_lmshift() with $p=1$
fit_trendshift_ar1() $42m+$ 4 $\sigma^{2},\phi,\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}$
fit_lmshift() $p(m+1)+1$ $\sigma^{2},\beta_{00},\ldots,\beta_{0m},\beta_{10},\ldots,\beta_{pm}$ polynomial [^1] Unlike fit_meanshift_norm(), fit_lmshift() function uses the lm() function from the stats, which provides flexibility at the expense of speed (fit_meanshift_norm() is faster). of degree $p$
fit_lmshift_ar1() $p(m+1)+1$ $\sigma^{2},\beta_{00},\ldots,\beta_{0m},\beta_{10},\ldots,\beta_{pm}$
fit_meanvar() $22m+$ 2 $\sigma_{0}^{2},\ldots,\sigma_{m}^{2},\mu_{0},\ldots,\mu_{m}$
fit_nhpp() $22m+$ 2 $\alpha_{0},\ldots,\alpha_{m},\beta_{0},\ldots,\beta_{m}$
@@ 5.3 Spotlight: Non-homogenous Poisson process model
Let $z_{w}(y)=\{t:y_{t}>w\}$ be a subset of the indices of the original times series $y$, for which the observation $y_{t}$ exceeds some threshold $w$, which by default is the empirical mean $\bar{y}$. This new time series $z_{w}(y)$ is called the exceedances of the original time series $y$. Following Suárez-Sierra, Coen, and Taimal (2023), we model the time series $z_{w}(y)$ as a non-homogenous Poisson process (NHPP), using a Weibull distribution parameterized by shape ($\alpha$) and scale ($\beta$). In a Bayesian setting, those parameters follow a Gamma distribution with their appropriate hyperparameters. Thus, we assume that there exists a set of parameters $\theta=(\alpha_{0}\ldots\alpha_{m},\beta_{0},\ldots,\beta_{m})$ for which the exceedances $z_{w}(y)$ of the original time series $y$ are modeled accurately as a non-homogenous Poisson process.
The best-fit parameters are determined by maximizing the log posterior function described below:
$\ln{g(\theta_{\tau}|y)}\propto\ln{L_{NHPP}(y|\theta_{\tau})}+\ln{g(\theta_{\tau})}\,.$
@@ 6 Penalized objective functions
We call the function $f$ a penalized objective function [^2] Occasionally, we may abuse terminology by referring to $f$ as a penalty function, but technically, $P_{f}$ is the penalty function. because it comes in the form
$f(\tau,M(y|\theta_{\tau}))=P_{f}(\tau,n)-2\ln{L_{M}(y|\theta_{\tau})}\,,$
where $P_{f}$ is a function that guards against overfitting by penalizing large changepoint sets—that is, it is a monotonically increasing function of $m$. The stats package (R Core Team 2024) provides generic methods for the well-known penalty functions AIC() and BIC(), which work seamlessly on tidychangepoint models because they all implement methods for logLik(). In addition, tidychangepoint provides generic functions and methods for four additional penalty functions: HQC(), MBIC(), MDL(), and BMDL(). For compatibility with changepoint we also provide the alias SIC() for BIC(). tidychangepoint supports all of the penalty functions supported by changepoint, with the exception of the “Asymptotic” and “CROPS” penalties, which are not fully generalizable.
@@
6.1 Penalized objective functions supported in tidychangepoint
6.1.1 Hannan-Quinn Information Criterion
The Hannan-Quinn information Criterion is a slight tweak on the BIC that involves an iterated logarithm (Hannan and Quinn 1979): $P_{HQC}(\tau,n)=2m\cdot\ln{\ln{n}}$.
6.1.2 Modified Bayesian Information Criterion
@@
$P_{MBIC}(\tau,n)=3m\ln{n}+\sum_{j=1}^{m+1}\ln{\frac{\ell_{j}}{n}}\,.$
6.1.3 Minimum Descriptive Length
@@
$P_{MDL}(\tau,n)=\frac{a(\theta_{\tau})}{2}\cdot\sum_{j=0}^{m}\log{\ell_{j}}+2\ln{m}+\sum_{j=2}^{m}\ln{\tau_{j}}+\left(2+b(\theta_{\tau})\right)\ln{n}\,.$
@@
6.1.4 Bayesian Minimum Descriptive Length
@@
$BMDL(\tau,NHPP(y|\hat{\theta}_{\tau}))=P_{MDL}(\tau)-2\ln{L_{NHPP}(y|\hat{\theta}_{\tau})}-2\ln{g(\hat{\theta}_{\tau})}\,.$
@@ 6.2 Extending tidychangepoint to include new penalty functions
 - attr(*, "df")= num 14
 - attr(*, "nobs")= int 95
 - attr(*, "tau")= int [1:4] 15 18 52 87
@@ 7 Results
In this section we document tidychangepoint’s correctness and computational costs.
7.1 Correctness
Since tidychangepoint relies on other packages to actually segment time series via their respective changepoint detection algorithms, the results returned by segment() are correct to the extent that the dependent results are correct.
However, recall that the model component of a tidycpt object is computed by our fit_*() model-fitting functions, after the segmenting algorithm has returned its results. We show below that these computations are accurate.
First, we match the model parameters returned by fit_trendshift() and fit_trendshift_ar1() with the results for the Central England Temperature data (CET) reported in rows 1–4 of Table 2 in Shi et al. (2022). Unfortunately, the CET data been revised since the publication of Shi et al. (2022), so the temperatures present in CET do not exactly match the temperatures used in Shi et al. (2022) for all years. Accordingly, the numbers in Table LABEL:tbl-shi-compare do not match exactly, but they are very close. However, when we perform this same analysis on the old data (not shown here), the numbers match exactly. [^3] Other results from Table 3 also match.
cpts <- c(170017391988)
ids <- time2tau(cpts, as_year(time(CET)))
trend_wn <- CET["/2020-01-01"|>
  fit_trendshift(tau = ids)
trend_ar1 <- CET["/2020-01-01"|>
  fit_trendshift_ar1(tau = ids)
Table 3: Comparison of results reported by Shi et al. (2022) to those reported by tidychangepoint. All discrepancies are caused by revisions to the underlying CET data used.
Model $\hat{\sigma}^{2}$ logLik BIC MDL $\hat{\phi}$
Shi et al. (2022) WN 0.291 -290.02 650.74 653.07 NA
trend_wn 0.290 -289.86 650.41 652.75 NA
Shi et al. (2022) AR(1) 0.290 -288.80 654.19 656.52 0.058
trend_ar1 0.290 -288.50 653.59 655.94 0.055
Second, in Table LABEL:tbl-param-compare we compare the estimated parameter values from the model component of a tidycpt object (which are fit by fit_meanvar()) to those from the segmenter component (which are fit by cpt.meanvar()) using a simulated data set.
tidycpt <- DataCPSim |>
  segment(method = "pelt"penalty = "BIC")
# Results shown in Table...
# tidycpt\$segmenter@param.est
# y\$model\$region_params
Table 4: Comparison of parameter values returned by fit_meanvar() ($\hat{\mu}_{tidycpt}$ and $\hat{\sigma}^{2}_{tidycpt}$) to those of changepoint::cpt.meanvar() ($\hat{\mu}_{cpt}$ and $\hat{\sigma}^{2}_{cpt}$). The discrepancies in the means are caused by tidychangepoint using intervals open on the right, as opposed to the left. Additional discrepancies in the variances are caused by changepoint’s use of the population variance.
region $\hat{\mu}_{tidycpt}$ $\hat{\sigma}^{2}_{tidycpt}$ $\hat{\mu}_{cpt}$ $\hat{\sigma}^{2}_{cpt}$
[1,547) 35.28 127.09 35.28 126.88
[547,822) 58.10 371.75 58.20 370.52
[822,972) 96.71 923.95 96.77 920.98
[972,1.1e+03) 155.85 2441.82 156.52 2405.97
The tiny differences in the means are because tidychangepoint uses intervals that are closed on the left and open on the right, whereas changepoint does the opposite. The differences in the variances are caused by the same issue and changepoint’s reporting of the population variance instead of the sample variance.
7.2 Benchmarking
Since segment() wraps underlying functionality from other packages, there is overhead created, both in terms of memory and speed. Table LABEL:tbl-bench shows the results of benchmarking segment() against changepoint::cpt.meanvar(), which is doing the underlying work in both cases. In this case, segment() uses about twice as much memory and is about 8 times slower than the underlying function. While that may seem like a lot, we’re talking about microseconds.
Table 5: Benchmarking
expression min median itr/sec mem_alloc gc/sec n_itr n_gc total_time
tidychangepoint 13.05ms 13.99ms 71.59874 75.7KB 0.0000 10 0 139.7ms
changepoint 1.45ms 1.56ms 595.09531 35.7KB 66.1217 9 1 15.1ms
As noted in Section 4.3, tidychangepoint provides two alternative functions to GA::gabin_Population() for creating an initial population for a genetic algorithm: build_gabin_population() and log_gabin_population(). Both perform some näive analysis to set the probability of an observation being selected uniformly at random to belong to the first generation. The former runs three quick algorithms (PELT, Binary Segmentation, and WBS) and sets the probability to three times the average size of the changepoint sets returned by those algorithms, divided by $n$. The latter sets the probability to $\log{n}/n$. Both functions can be thought of as “informed” by the data, albeit in different ways. Recall that GA::gabin_Population() always uses the probability $1/2$, and can thus be considered “uninformed”.
In Figure 11, we compare the performance of a genetic algorithm on monthly retail data from Australia (O’Hara-Wild et al. 2022), using these three methods of forming the initial population. In this test, the maximum number of iterations was set to 1000, with a consecutive run of 100 generations with no improvement stopping the algorithm. We see in Figure 11 that both of the “informed” initial populations result in improved performance. The recovered changepoint sets score better values of the penalized objective function, and both algorithms terminated well before they reached the maximum number of generations. Conversely, the algorithm using the default initial population was still going after 1000 generations. In fact, it only took 35 generations for the BIC scores returned by the genetic algorithm whose initial population was set by log_gabin_population() to beat the best score returned by the same algorithm using the default initial population after 1000 generations.
Figure 11: Comparison the evolution of the best candidate changepoint sets for genetic algorithms with different initial populations. The ‘runif’ series begins by selecting observations to belong to the candidate changepoint set uniformly at random, with probability 0.5. On the other hand, the ‘build’ and ‘logn’ series also selects candidate changepoints uniformly at random, but with informed probabilities.
Of course, it is possible that by casting a wider net, the default initial population will catch changepoint sets that were randomly excluded by the “informed” initial populations. But these results held up under repeated testing, and they suggest a promising line of attack for future work.